OpenStudy (anonymous):

can you help me to solve this integral and explain me?? I can t undertand :/ (x^5 + 2x^2)/ (x^6 + 2x^3+ 5 ) dx

OpenStudy (caozeyuan):

ok, seems not that hard

OpenStudy (caozeyuan):

\[\int\limits \frac{ x^5+2x ^{2} }{ x ^{6}+2x ^{3}+5}dx\]

OpenStudy (caozeyuan):

\[\int\limits \frac{ f \prime(x) }{ f(x) }dx=\ln f(x)\]

OpenStudy (caozeyuan):

wait no, his is way harder han I thought

OpenStudy (caozeyuan):

can you check that it is indeed 2x squared, not x squared

OpenStudy (caozeyuan):

@almavera93

OpenStudy (gorv):

x^6 + 2x^3+ 5= we can write it as \[(x^3)^2+2*x^3+1+4\]

OpenStudy (gorv):

\[(x^3+1)^2+4\]

OpenStudy (gorv):

\[\int\limits \frac{ x^5+2x^2 }{ (x^3+1)^2+4 } * dx\]

OpenStudy (gorv):

\[\int\limits \frac{ x^5+x^2+x^2 }{ (x^3+1)^2+4 }*dx\]

OpenStudy (gorv):

\[\int\limits \frac{ (x^3+1)*x^2+x^2 }{ (x^3+1)^2+4 }*dx\]

OpenStudy (gorv):

\[\int\limits \frac{ (x^3+1)*x^2*dx }{ (x^3+1)^2+4 } + \int\limits\frac{ x^2*dx }{ (x^3+1)^2+4} *dx\]

OpenStudy (gorv):

now let x^3+1=t on differentiating 3x^2*dx=dt x^2dx=dt/3 put these 2 in our integral

OpenStudy (gorv):

\[\int\limits \frac{ t*\frac{ dt }{ 3 } }{ t^2+4 }+ \int\limits \frac{ \frac{ dt }{ 3 }}{ t^2+4 }\]

OpenStudy (gorv):

\[\frac{ 1 }{ 3 } \int\limits \frac{ t*dt }{ t^2+4 }+\frac{ 1 }{ 3 }* \int\limits \frac{ dt }{ t^2+4}\]

OpenStudy (gorv):

now its simple try it now

OpenStudy (anonymous):

@gorv 0oh thank you, that way it is easier